A bag contains 25-paisa and 50-paise coins whose total value is rs.30.If the number of 25-paisa coins is four times that of 50-paisa coins.Find the number of each type of coins.
Answer by mananth(16946)

50 paise coins be x
25 paise coins will be 4x
Total vaue of the coins = RS.30 = 300 PAISE
50x+25*4x= 300
50x+100x=300
150x=300
x=2
2 50 paise coins
25 paise coins = 4x= 4*2=8
2, 50 paise coins & 8, 25 paise coins